OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (anonymous):

(y-y0)=m(x-x0) (y-(-3))=-6(x-(-2)) y+3=-6x-12 y=-6x-15

OpenStudy (anonymous):

it is true?